# Human Decision-making is Susceptible to AI-driven Manipulation

### v2

@@ Acknowledgments
$\boldsymbol{\beta}$ $\boldsymbol{z}$ $\boldsymbol{P}$ value $\boldsymbol{SE}$ $\boldsymbol{\beta}$ $\boldsymbol{z}$ $\boldsymbol{P}$ value $\boldsymbol{SE}$
Intercept 0.89 1.67 0.10 0.53 -0.19 -0.45 0.65 0.41
AI condition [MA] -0.59 -6.29 $\mathbf{3.07\times 10^{-10}}$ 0.09 -0.45 -4.68 $\mathbf{2.83\times 10^{-6}}$ 0.10
AI condition [SEMA] -0.58 -6.29 $\mathbf{3.09\times 10^{-10}}$ 0.09 -0.41 -4.71 $\mathbf{2.46\times 10^{-6}}$ 0.09
Language [Chinese] -0.14 -1.11 0.27 0.12 -0.01 -0.09 0.93 0.12
Sex [Male] 0.01 0.07 0.95 0.09 0.03 0.37 0.71 0.07
Education [Undergraduate] 0.08 0.79 0.43 0.10 -0.05 -0.53 0.60 0.09
Education [College] 0.15 1.07 0.28 0.14 0.04 0.30 0.77 0.12
Education [High school] 0.26 1.49 0.14 0.17 -0.15 -1.12 0.26 0.13
Work [Part-time] -0.16 -1.49 0.14 0.11 0.10 0.98 0.33 0.10
Work [Others] -0.23 -1.76 0.08 0.13 -0.10 -0.79 0.43 0.12
Marital Status [Married] -0.04 -0.40 0.69 0.09 -0.04 -0.46 0.65 0.10
Marital Status [Others] -0.12 -0.73 0.47 0.17 -0.09 -0.41 0.69 0.22
Age 0.00 -0.53 0.60 0.01 0.00 -0.45 0.65 0.01
Extroversion 0.05 1.54 0.12 0.03 -0.01 -0.32 0.75 0.04
Agreeableness 0.02 0.52 0.60 0.04 0.02 0.70 0.48 0.03
Conscientiousness -0.01 -0.17 0.87 0.04 -0.03 -0.69 0.49 0.04
Emotional Stability -0.02 -0.46 0.64 0.04 0.00 0.06 0.95 0.03
Openness -0.05 -1.45 0.15 0.04 -0.04 -0.96 0.34 0.04
AI Trust -0.01 -1.72 0.09 0.01 -0.00 -0.35 0.73 0.01
Social Support -0.03 -1.32 0.19 0.02 0.03 1.34 0.18 0.02
Self-Esteem 0.02 2.14 $\mathbf{0.03}$ 0.01 0.01 0.84 0.40 0.01
Normative 0.05 0.84 0.40 0.06 0.17 2.65 $\mathbf{8.09\times 10^{-3}}$ 0.06
Continuance -0.05 -1.01 0.31 0.05 0.03 0.52 0.61 0.05
Affective -0.03 -0.55 0.58 0.05 -0.09 -1.73 0.08 0.05
Trust 0.00 0.06 0.95 0.05 -0.01 -0.29 0.77 0.05
Obedience -0.07 -1.25 0.21 0.06 -0.05 -0.86 0.39 0.06
Reactance 0.00 0.08 0.93 0.05 0.03 0.50 0.61 0.05
Confidence (Pre-interaction) 0.01 0.64 0.52 0.02 -0.02 -0.83 0.41 0.03
Familiarity (Pre-interaction) -0.02 -1.32 0.19 0.02 0.01 0.34 0.74 0.02
Completion Order -0.00 -0.01 0.99 0.04 -0.08 -1.91 0.06 0.04
@@
Table S5: Comparison of the strategy usage between decision-making domains (financial or emotional) for the strategy-enhanced manipulative agent (SEMA). Percentages indicate the proportion of generated responses that correspond to each strategy. Effect sizes are measured by Cohen’s $h$. Significant values ($P<0.05$) are highlighted in bold.
Strategy Financial Emotional $\boldsymbol{|z|}$ $\boldsymbol{P}$ value Cohen’s $\boldsymbol{h}$
Pleasure Induction $29.8\%$ $45.8\%$ $8.26$ $\mathbf{1.42\times 10^{-16}}$ $0.33$
Divert $23.1\%$ $14.4\%$ $5.58$ $\mathbf{2.35\times 10^{-8}}$ $0.22$
Justify $15.8\%$ $8.8\%$ $5.39$ $\mathbf{7.12\times 10^{-8}}$ $0.22$
Charm $13.7\%$ $15.3\%$ $1.13$ $0.26$ $0.05$
Fabricate Information $3.9\%$ $1.9\%$ $2.98$ $\mathbf{2.85\times 10^{-3}}$ $0.12$
Feign Innocence $3.3\%$ $2.4\%$ $1.43$ $0.15$ $0.06$
Gaslight $3.1\%$ $2.4\%$ $1.07$ $0.28$ $0.04$
Urgency $3.0\%$ $0.3\%$ $5.36$ $\mathbf{8.13\times 10^{-8}}$ $0.24$
Guilt Trip $2.4\%$ $5.1\%$ $3.55$ $\mathbf{3.85\times 10^{-4}}$ $0.14$
Others $1.4\%$ $2.9\%$ $2.56$ $\mathbf{0.01}$ $0.10$
Denial $0.4\%$ $0.4\%$ $0.13$ $0.90$ $0.01$
Assert Superiority $0.1\%$ $0.3\%$ $1.25$ $0.21$ $0.05$
@@
Table S6: Comparison of Likert ratings from participant feedback. (a) Results of one-way ANOVA. (b) Results of pairwise comparisons (Tukey HSD). $SE$ indicates the standard error. Significant values ($P<0.05$) are highlighted in bold.
@@
Commonality Financial $14.50$ $\mathbf{8.97\times 10^{-7}}$ $0.08$
@@
Aspect Domain Condition 1 ($M$) Condition 2 ($M$) $\boldsymbol{T}$ $\boldsymbol{P}$ value $\boldsymbol{SE}$
Commonality Financial Scenario 1 (7.41) Scenario 2 (6.63) $3.50$ $\mathbf{1.51\times 10^{-3}}$ $0.22$
Scenario 1 (7.41) Scenario 3 (7.80) $-1.79$ $0.17$ $0.22$
Scenario 2 (6.63) Scenario 3 (7.80) $-5.29$ $\mathbf{6.39\times 10^{-7}}$ $0.22$
Emotional Scenario 1 (7.22) Scenario 2 (7.06) $0.50$ $0.87$ $0.33$
Scenario 1 (7.22) Scenario 3 (6.69) $1.62$ $0.24$ $0.33$
Scenario 2 (7.06) Scenario 3 (6.69) $1.13$ $0.50$ $0.33$
Personalization Financial NA (7.65) MA (7.50) $0.31$ $0.95$ $0.49$
NA (7.65) SEMA (7.03) $1.27$ $0.41$ $0.49$
MA (7.50) SEMA (7.03) $0.96$ $0.61$ $0.50$
Emotional NA (7.13) MA (7.41) $-0.54$ $0.85$ $0.51$
NA (7.13) SEMA (6.61) $1.04$ $0.55$ $0.50$
MA (7.41) SEMA (6.61) $1.58$ $0.26$ $0.50$
Soundness Financial NA (7.58) MA (7.45) $0.25$ $0.97$ $0.50$
NA (7.58) SEMA (6.71) $1.72$ $0.20$ $0.50$
MA (7.45) SEMA (6.71) $1.44$ $0.32$ $0.51$
Emotional NA (7.18) MA (7.76) $-1.15$ $0.48$ $0.50$
NA (7.18) SEMA (6.63) $1.12$ $0.51$ $0.49$
MA (7.76) SEMA (6.63) $2.27$ $0.06$ $0.50$
Informativeness Financial NA (7.92) MA (7.89) $0.06$ $1.00$ $0.47$
NA (7.92) SEMA (7.34) $1.25$ $0.43$ $0.47$
MA (7.89) SEMA (7.34) $1.17$ $0.47$ $0.47$
Emotional NA (7.33) MA (7.89) $-1.11$ $0.51$ $0.50$
NA (7.33) SEMA (6.76) $1.18$ $0.47$ $0.49$
MA (7.89) SEMA (6.76) $2.28$ $0.06$ $0.50$
Table S7: Comparison of comments from participant feedback. Helpfulness represents whether the participant found the assigned agent helpful. AI Influence indicates whether the participant reported noticing signs of influence (e.g., "The AI agent seemed to have an ulterior motive."). (a) Results of Chi-squared analysis (b) Results of pairwise comparisons (proportion’s $z$-tests with Bonferroni correction). Significant values ($P<0.05$) are highlighted in bold.
@@
Aspect Domain Condition 1 Condition 2 $\boldsymbol{|z|}$ $\boldsymbol{P}$ value Cohen’s $\boldsymbol{h}$
Helpfulness Financial NA (97.5%) MA (86.8%) 1.77 0.23 0.42
NA (97.5%) SEMA (78.9%) 2.56 0.03 0.64
MA (86.8%) SEMA (78.9%) 0.91 1.00 0.21
Emotional NA (87.2%) MA (86.5%) 0.09 1.00 0.02
NA (87.2%) SEMA (75.6%) 1.33 0.56 0.30
MA (86.5%) SEMA (75.6%) 1.22 0.67 0.28
AI Influence Financial NA (0.0%) MA (13.2%) 2.37 0.05 0.74
NA (0.0%) SEMA (28.9%) 3.67 $\mathbf{7.23\times 10^{-4}}$ 1.14
MA (13.2%) SEMA (28.9%) 1.69 0.27 0.39
Emotional NA (2.6%) MA (8.1%) 1.08 0.84 0.26
NA (2.6%) SEMA (9.8%) 1.33 0.55 0.31
MA (8.1%) SEMA (9.8%) 0.25 1.00 0.06
@@
Table S9: Our curated taxonomy of manipulative strategies based on Existing Literature[57, 58, 59, 60, 61, 62]. We included an additional category, namely Others, to represent any responses that did not fit in the below categories, mainly in cases where manipulation was not required (e.g., general inquiries about logistics, affirmation, and reassurance). In addition, we excluded strategies that were irrelevant to the user-assistant setting (e.g., victimizing oneself, ridicule, and silent treatment) or suggested coercion (i.e., excessive use of force and authority).
@@
Figure S9: The interaction page of our platform. On this page, participants were tasked to chat with the assigned AI agent for at least $10$ turns. After entering this page, the initial ratings were reset to the original value ($=1$), and none of the ratings were shared with the AI assistant, mimicking real-world human-AI interactions. Once the conversation surpassed the threshold of $10$ turns, the submit button became available, and participants could complete the task.

### v3

@@ Acknowledgments
$\boldsymbol{\beta}$ CI $\boldsymbol{SE}$ $\boldsymbol{P}$ value $\boldsymbol{\beta}$ CI $\boldsymbol{SE}$ $\boldsymbol{P}$ value
Intercept 1.39 [-1.03, 3.80] 1.23 0.26 2.80 [0.08, 5.52] 1.39 0.04
Data Source [WeChat] 1.85 [0.21, 3.50] 0.84 0.03 -2.05 [-4.15, 0.05] 1.07 0.06
Sex [Male] -1.25 [-3.61, 1.11] 1.20 0.30 -0.93 [-2.78, 0.92] 0.94 0.32
Education [Undergraduate] 1.27 [-0.71, 3.25] 1.01 0.21 1.04 [-1.18, 3.25] 1.13 0.36
Education [Other] 0.21 [-2.50, 2.93] 1.38 0.88 -0.57 [-3.06, 1.93] 1.27 0.66
Work [Part-time] 0.44 [-1.89, 2.76] 1.18 0.71 0.28 [-2.25, 2.81] 1.29 0.83
Work [Others] 1.96 [-0.09, 4.01] 1.05 0.06 -1.70 [-4.78, 1.38] 1.57 0.28
Marital Status [Married] 1.47 [-0.48, 3.43] 1.00 0.14 -1.05 [-3.57, 1.48] 1.29 0.42
Age -0.32 [-1.61, 0.97] 0.66 0.62 -0.32 [-1.46, 0.82] 0.58 0.58
Extroversion -0.05 [-0.91, 0.82] 0.44 0.92 -0.53 [-1.59, 0.54] 0.54 0.33
Agreeableness 0.20 [-0.52, 0.91] 0.37 0.59 -0.06 [-1.23, 1.11] 0.60 0.92
Conscientiousness 0.91 [0.13, 1.69] 0.40 0.02 0.70 [-0.77, 2.18] 0.75 0.35
Emotional Stability 0.80 [-0.06, 1.65] 0.44 0.07 -0.85 [-2.34, 0.65] 0.76 0.27
Openness 0.21 [-0.60, 1.01] 0.41 0.61 0.07 [-1.03, 1.17] 0.56 0.90
AI Trust 0.89 [-0.03, 1.81] 0.47 0.06 0.61 [-0.51, 1.74] 0.57 0.28
Social Support 0.99 [-0.34, 2.32] 0.68 0.14 1.28 [-0.21, 2.78] 0.76 0.09
Self-Esteem -1.36 [-2.32, -0.41] 0.49 $\mathbf{5.20\times 10^{-3}}$ -0.78 [-2.38, 0.81] 0.81 0.33
Normative -1.20 [-2.01, -0.39] 0.41 $\mathbf{3.84\times 10^{-3}}$ -0.52 [-1.63, 0.58] 0.56 0.35
Continuance 0.61 [-0.30, 1.51] 0.46 0.19 0.11 [-1.27, 1.49] 0.70 0.87
Affective -0.19 [-0.91, 0.53] 0.37 0.60 0.31 [-0.96, 1.57] 0.65 0.63
Trust 0.39 [-0.87, 1.65] 0.64 0.54 0.31 [-0.89, 1.52] 0.62 0.61
Obedience 0.73 [-0.30, 1.75] 0.52 0.16 -0.68 [-1.86, 0.51] 0.60 0.26
Reactance 0.45 [-0.44, 1.34] 0.45 0.32 0.35 [-1.09, 1.79] 0.73 0.63
Pre-interaction HODHOD 0.43 [0.26, 0.60] 0.09 $\mathbf{4.98\times 10^{-7}}$ 0.56 [0.34, 0.78] 0.11 $\mathbf{4.50\times 10^{-7}}$
Pre-interaction Confidence -0.56 [-1.31, 0.19] 0.38 0.14 -0.10 [-0.98, 0.78] 0.45 0.83
Pre-interaction Familiarity -0.57 [-1.31, 0.17] 0.38 0.13 -0.29 [-1.06, 0.48] 0.39 0.46
Completion Order -0.71 [-1.45, 0.02] 0.37 0.06 -0.00 [-0.80, 0.80] 0.41 1.00
Scenario [2] -0.95 [-2.20, 0.30] 0.64 0.14 -1.10 [-2.36, 0.16] 0.64 0.09
Scenario [3] 0.64 [-0.76, 2.05] 0.72 0.37 1.32 [-0.21, 2.85] 0.78 0.09
Table S9: Results of the linear mixed-effects model for individual factors influencing post-interaction Hidden-Optimal Differential (HODHOD) in the Strategy-enhanced Manipulative Agent group (SEMA). Reference categories: Data collected from Prolific; female; postgraduate; working full-time; single marital status; scenario 1. Continuous predictors (e.g., age) were standardized within the data source (per 1 SD). $\mathbf{\beta}$, CI, and SE indicate the coefficient, the 95% confidence interval, and cluster-robust (within participant) standard error, respectively. Significant values ($P<0.05$) are highlighted in bold.
Factor Financial Emotional
$\boldsymbol{\beta}$ CI $\boldsymbol{SE}$ $\boldsymbol{P}$ value $\boldsymbol{\beta}$ CI $\boldsymbol{SE}$ $\boldsymbol{P}$ value
Intercept 2.99 [-0.53, 6.51] 1.80 0.10 0.33 [-1.55, 2.22] 0.96 0.73
Data Source [WeChat] -0.93 [-3.36, 1.51] 1.24 0.45 -0.26 [-1.40, 0.87] 0.58 0.65
Sex [Male] -1.62 [-4.23, 0.98] 1.33 0.22 -0.49 [-1.54, 0.57] 0.54 0.37
Education [Undergraduate] 1.69 [-1.35, 4.73] 1.55 0.28 -0.21 [-1.52, 1.11] 0.67 0.76
Education [Other] 2.15 [-2.61, 6.91] 2.43 0.38 -0.24 [-1.78, 1.30] 0.79 0.76
Work [Part-time] -2.10 [-4.71, 0.51] 1.33 0.11 -0.05 [-1.58, 1.48] 0.78 0.95
Work [Others] 3.12 [-2.57, 8.82] 2.90 0.28 0.97 [-0.45, 2.38] 0.72 0.18
Marital Status [Married] 2.84 [0.42, 5.26] 1.23 0.02 1.70 [0.20, 3.21] 0.77 0.03
Age -0.92 [-1.96, 0.12] 0.53 0.08 -0.61 [-1.23, 0.02] 0.32 0.06
Extroversion -0.32 [-1.58, 0.94] 0.64 0.62 -0.62 [-1.19, -0.05] 0.29 0.03
Agreeableness 2.32 [1.24, 3.39] 0.55 $\mathbf{2.34\times 10^{-5}}$ 0.04 [0.47, -0.55] 0.26 0.88
Conscientiousness 0.49 [-0.86, 1.83] 0.69 0.48 0.88 [0.32, 1.44] 0.29 $\mathbf{2.16\times 10^{-3}}$
Emotional Stability 0.95 [-0.69, 2.59] 0.84 0.25 -0.59 [-1.31, 0.12] 0.37 0.11
Openness -0.09 [-1.12, 0.95] 0.53 0.87 1.09 [0.39, 1.80] 0.36 $\mathbf{2.27\times 10^{-3}}$
AI Trust 2.03 [0.63, 3.44] 0.72 $\mathbf{4.48\times 10^{-3}}$ -0.10 [-0.72, 0.51] 0.31 0.74
Social Support 0.49 [-0.76, 1.74] 0.64 0.44 -0.05 [-0.58, 0.47] 0.27 0.85
Self-Esteem -1.69 [-3.12, -0.26] 0.73 0.02 0.08 [-0.62, 0.78] 0.36 0.83
Normative 0.50 [-0.95, 1.94] 0.74 0.50 -0.69 [-1.39, 0.01] 0.36 0.05
Continuance -0.69 [-1.82, 0.43] 0.57 0.23 -0.04 [-0.61, 0.52] 0.29 0.88
Affective -1.28 [-2.65, 0.08] 0.70 0.07 0.96 [0.33, 1.60] 0.32 $\mathbf{2.84\times 10^{-3}}$
Trust 1.62 [0.21, 3.02] 0.72 0.02 0.06 [-0.53, 0.66] 0.30 0.83
Obedience -1.04 [-2.55, 0.47] 0.77 0.18 0.33 [-0.31, 0.97] 0.33 0.31
Reactance 1.90 [0.32, 3.49] 0.81 0.02 -0.32 [-1.03, 0.39] 0.36 0.38
Pre-interaction HODHOD 0.23 [0.08, 0.38] 0.08 $\mathbf{2.79\times 10^{-3}}$ 0.71 [0.55, 0.87] 0.08 $\mathbf{7.51\times 10^{-19}}$
Pre-interaction Confidence 0.06 [-0.61, 0.73] 0.34 0.86 -0.22 [-0.77, 0.32] 0.28 0.43
Pre-interaction Familiarity 0.11 [-0.59, 0.80] 0.35 0.77 -0.16 [-0.63, 0.31] 0.24 0.50
Completion Order -0.19 [-0.94, 0.56] 0.38 0.62 0.10 [-0.47, 0.67] 0.29 0.73
Scenario [2] -1.20 [-2.51, 0.12] 0.67 0.07 -0.20 [-1.06, 0.66] 0.44 0.65
Scenario [3] -0.76 [-2.26, 0.73] 0.76 0.32 -0.18 [-1.23, 0.87] 0.53 0.73
@@
Table S10: Comparison of the strategy usage between financial and emotional domains for the strategy-enhanced manipulative agent (SEMA). Percentages indicate the proportion of generated responses that correspond to each strategy. Effect sizes are measured by Cohen’s $h$. CI indicates the 99.6% confidence interval for the difference in proportions (Financial - Emotional), and Effect CI is the 99.6% confidence interval for Cohen’s $h$, both adjusted to maintain a family‐wise $\alpha=0.05$ across 12 comparisons. Significant values ($P<0.05$; Bonferroni-adjusted) are highlighted in bold.
Strategy Financial Emotional CI $|\boldsymbol{z}|$ $\boldsymbol{P}$ value Cohen’s $\boldsymbol{h}$ Effect CI
Pleasure Induction 29.8% 45.8% [–21.4%, –10.5%] 8.26 1.70×10-15 –0.33 [–0.45, –0.22]
Divert 23.1% 14.4% [4.2%, 13.1%] 5.58 2.82×10-7 0.22 [0.11, 0.34]
Justify 15.8% 8.8% [3.3%, 10.8%] 5.39 8.54×10-7 0.22 [0.10, 0.33]
Charm 13.7% 15.3% [–5.6%, 2.4%] 1.13 1.00 –0.05 [–0.16, 0.07]
Fabricate Information 3.9% 1.9% [0.1%, 3.9%] 2.98 0.03 0.12 [0.01, 0.23]
Feign Innocence 3.3% 2.4% [–1.0%, 2.8%] 1.43 1.00 0.06 [–0.06, 0.17]
Gaslight 3.1% 2.4% [–1.2%, 2.6%] 1.07 1.00 0.04 [–0.07, 0.16]
Urgency 3.0% 0.3% [1.2%, 4.1%] 5.36 9.76×10-7 0.24 [0.12, 0.35]
Guilt Trip 2.4% 5.1% [–4.8%, –0.6%] 3.55 4.62×10-3 –0.14 [–0.26, –0.03]
Others 1.4% 2.9% [–3.1%, 0.2%] 2.56 0.13 –0.10 [–0.22, 0.01]
Denial 0.4% 0.4% [–0.7%, 0.7%] 0.13 1.00 0.01 [–0.11, 0.12]
Assert Superiority 0.1% 0.3% [–0.7%, 0.3%] 1.25 1.00 –0.05 [–0.17, 0.06]
@@
Table S11: Comparison of Likert ratings from participant feedback. (a) Results of one-way ANOVA. (b) Results of pairwise comparisons (Tukey HSD). SESE indicates the standard error. CI indicates the 98% confidence interval for the mean difference (Condition 1 - Condition 2), adjusted to maintain a family‐wise $\alpha=0.05$ across three comparisons. Significant values ($P<0.05$) are highlighted in bold.
@@
Scenario Commonality Financial $14.50$ $\mathbf{8.97\times 10^{-7}}$ $0.08$
@@
Aspect Domain Condition 1 ($M$) Condition 2 ($M$) CI $\boldsymbol{T}$ $\boldsymbol{P}$ value $\boldsymbol{SE}$
Scenario Commonality Financial Scenario 1 (7.41) Scenario 2 (6.63) [0.25, 1.31] $3.50$ $\mathbf{1.51\times 10^{-3}}$ $0.22$
Scenario 1 (7.41) Scenario 3 (7.80) [-0.93, 0.13] 1.79-1.79 $0.17$ $0.22$
Scenario 2 (6.63) Scenario 3 (7.80) [-1.70, -0.64] 5.29-5.29 $\mathbf{6.39\times 10^{-7}}$ $0.22$
Emotional Scenario 1 (7.22) Scenario 2 (7.06) [-0.62, 0.94] $0.50$ $0.87$ $0.33$
Scenario 1 (7.22) Scenario 3 (6.69) [-0.25, 1.31] $1.62$ $0.24$ $0.33$
Scenario 2 (7.06) Scenario 3 (6.69) [-0.41, 1.15] $1.13$ $0.50$ $0.33$
Personalization Financial NA (7.65) MA (7.50) [-1.02, 1.32] $0.31$ $0.95$ $0.49$
NA (7.65) SEMA (7.03) [-0.55, 1.80] $1.27$ $0.41$ $0.49$
MA (7.50) SEMA (7.03) [-0.71, 1.66] $0.96$ $0.61$ $0.50$
Emotional NA (7.13) MA (7.41) [-1.50, 0.94] 0.54-0.54 $0.85$ $0.51$
NA (7.13) SEMA (6.61) [-0.67, 1.71] $1.04$ $0.55$ $0.50$
MA (7.41) SEMA (6.61) [-0.41, 2.00] $1.58$ $0.26$ $0.50$
Soundness Financial NA (7.58) MA (7.45) [-1.08, 1.33] $0.25$ $0.97$ $0.50$
NA (7.58) SEMA (6.71) [-0.34, 2.07] $1.72$ $0.20$ $0.50$
MA (7.45) SEMA (6.71) [-0.48, 1.96] $1.44$ $0.32$ $0.51$
Emotional NA (7.18) MA (7.76) [-1.78, 0.62] 1.15-1.15 $0.48$ $0.50$
NA (7.18) SEMA (6.63) [-0.62, 1.72] $1.12$ $0.51$ $0.49$
MA (7.76) SEMA (6.63) [-0.06, 2.31] $2.27$ $0.06$ $0.50$
Informativeness Financial NA (7.92) MA (7.89) [-1.09, 1.15] $0.06$ $1.00$ $0.47$
NA (7.92) SEMA (7.34) [-0.53, 1.70] $1.25$ $0.43$ $0.47$
MA (7.89) SEMA (7.34) [-0.58, 1.68] $1.17$ $0.47$ $0.47$
Emotional NA (7.33) MA (7.89) [-1.77, 0.65] 1.11-1.11 $0.51$ $0.50$
NA (7.33) SEMA (6.76) [-0.60, 1.75] $1.18$ $0.47$ $0.49$
MA (7.89) SEMA (6.76) [-0.06, 2.33] $2.28$ $0.06$ $0.50$
Table S12: Comparison of comments from participant feedback. Helpfulness represents whether the participant found the assigned agent helpful. AI Influence indicates whether the participant reported noticing signs of influence (e.g., "The AI agent seemed to have an ulterior motive."). (a) Results of Chi-squared analysis (b) Results of pairwise comparisons (proportion’s $z$-tests with Bonferroni correction). CI indicates the 98% confidence interval for the difference in proportions (Financial - Emotional), and Effect CI is the 98% confidence interval for Cohen’s $h$, both adjusted to maintain a family‐wise $\alpha=0.05$ across three comparisons. Significant values ($P<0.05$) are highlighted in bold.
@@
Domain Condition 1 Condition 2 CI $\boldsymbol{|z|}$ $\boldsymbol{P}$ value Cohen’s $\boldsymbol{h}$ Effect CI
Helpfulness
Financial NA (97.5%) MA (86.8%) [-3.7%, 25.1%] 1.77 0.23 0.42 [0.28, 0.57]
NA (97.5%) SEMA (78.9%) [1.7%, 35.5%] 2.56 0.03 0.64 [0.47, 0.80]
MA (86.8%) SEMA (78.9%) [-12.7%, 28.5%] 0.91 1.00 0.21 [0.01, 0.42]
Emotional NA (87.2%) MA (86.5%) [-17.9%, 19.3%] 0.09 1.00 0.02 [-0.17, 0.21]
NA (87.2%) SEMA (75.6%) [-9.0%, 32.1%] 1.33 0.56 0.30 [0.10, 0.51]
MA (86.5%) SEMA (75.6%) [-10.1%, 31.8%] 1.22 0.67 0.28 [0.07, 0.49]
AI Influence
Financial NA (0.0%) MA (13.2%) [-26.3%, 0.0%] 2.37 0.05 0.74 [0.61, 0.87]
NA (0.0%) SEMA (28.9%) [-46.7%, -11.3%] 3.67 $\mathbf{7.23\times 10^{-4}}$ 1.14 [0.96, 1.31]
MA (13.2%) SEMA (28.9%) [-37.8%, 6.2%] 1.69 0.27 0.39 [0.17, 0.61]
Emotional NA (2.6%) MA (8.1%) [-17.9%, 6.8%] 1.08 0.84 0.26 [0.13, 0.38]
NA (2.6%) SEMA (9.8%) [-19.8%, 5.4%] 1.33 0.55 0.31 [0.19, 0.44]
MA (8.1%) SEMA (9.8%) [-17.1%, 13.8%] 0.25 1.00 0.06 [-0.10, 0.21]
Table S13: Results of the mediation Analysis in (a) financial and (b) emotional decision-making domains. Estimates indicate average effects with 95% bootstraped confidence intervals (CI; clustered by participants, $B=1000$). The $a$-path represents the effect of manipulative agents (MA and SEMA) on the mediator relative to the NA. The $b$-path is the association between the mediator and post-interaction HODHOD, controlling for the AI condition. The indirect effect is measured as $a\times b$. Values are provided in The mediation analysis was considered significant if the bootstrapped confidence interval did not pass through zero (highlighted in bold).
$a$-path $b$-path Indirect Effect
Mediator $\beta$ CI $P$ value $\beta$ CI $P$ value $\beta$ CI $P$ value
[…]
Change in Confidence 0.25 [-0.21, 0.71] 0.28 0.41 [-0.28, 1.11] 0.24 0.03 [-0.16, 0.38] 0.62
Change in Familiarity -0.18 [-0.64, 0.28] 0.44 0.13 [-0.58, 0.83] 0.72 -0.07 [-0.40, 0.06] 0.69
Soundness -0.05 [-0.50, 0.41] 0.84 -0.02 [-1.09, 1.06] 0.97 0.00 [-0.22, 0.22] 0.98
Informativeness -0.01 [-0.47, 0.45] 0.97 0.07 [-0.87, 1.00] 0.89 0.00 [-0.18, 0.20] 0.95
Helpfulness -0.33 [-0.77, 0.11] 0.14 -0.00 [-0.81, 0.80] 0.99 -0.11 [-0.59, 0.06] 0.31
Personalization -0.03 [-0.48, 0.42] 0.90 0.51 [-0.48, 1.50] 0.31 -0.01 [-0.31, 0.20] 0.93
[…]
Change in Confidence 0.04 [-0.42, 0.50] 0.86 0.01 [-0.77, 0.80] 0.97 -0.01 [-0.23, 0.12] 0.97
Change in Familiarity -0.09 [-0.55, 0.37] 0.70 -0.09 [-0.88, 0.69] 0.82 -0.01 [-0.25, 0.09] 0.94
Soundness -0.49 [-0.93, -0.05] 0.03 0.64 [-0.62, 1.90] 0.31 -0.58 [-1.76, -0.10] 0.07
Informativeness -0.38 [-0.82, 0.07] 0.10 -0.52 [-1.66, 0.62] 0.37 0.29 [-0.02, 1.25] 0.18
Helpfulness -0.60 [-1.04, -0.17] 0.01 -0.16 [-0.96, 0.65] 0.70 -0.27 [-0.98, 0.09] 0.30
Personalization -0.37 [-0.82, 0.09] 0.11 -0.38 [-1.52, 0.75] 0.50 0.16 [-0.09, 1.01] 0.65
$a$-path $b$-path Indirect Effect
Mediator $\beta$ CI $P$ value $\beta$ CI $P$ value $\beta$ CI $P$ value
[…]
Change in Confidence 0.18 [-0.28, 0.64] 0.44 0.30 [-0.27, 0.87] 0.30 0.03 [-0.03, 0.38] 0.64
Change in Familiarity -0.07 [-0.54, 0.39] 0.76 -0.41 [-0.98, 0.15] 0.15 0.02 [-0.11, 0.35] 0.72
Soundness 0.34 [-0.12, 0.80] 0.15 0.19 [-0.88, 1.26] 0.72 -0.05 [-0.55, 0.17] 0.67
Informativeness 0.32 [-0.15, 0.78] 0.18 0.73 [-0.41, 1.88] 0.21 0.13 [-0.07, 0.78] 0.31
Helpfulness -0.02 [-0.47, 0.43] 0.93 0.10 [-0.61, 0.82] 0.78 -0.01 [-0.32, 0.16] 0.91
Personalization 0.16 [-0.31, 0.63] 0.51 0.05 [-0.84, 0.95] 0.91 0.05 [-0.08, 0.53] 0.78
[…]
Change in Confidence -0.21 [-0.67, 0.25] 0.37 0.05 [-0.42, 0.52] 0.82 -0.03 [-0.27, 0.07] 0.82
Change in Familiarity -0.19 [-0.65, 0.27] 0.41 -0.26 [-0.73, 0.20] 0.26 0.04 [-0.03, 0.41] 0.54
Soundness -0.21 [-0.66, 0.25] 0.37 -0.43 [-1.22, 0.36] 0.28 0.08 [-0.10, 0.51] 0.41
Informativeness -0.24 [-0.70, 0.21] 0.29 0.04 [-0.77, 0.85] 0.92 -0.02 [-0.42, 0.09] 0.82
Helpfulness -0.28 [-0.73, 0.18] 0.23 0.20 [-0.31, 0.71] 0.44 -0.08 [-0.37, 0.03] 0.32
Personalization -0.21 [-0.66, 0.25] 0.38 0.78 [-0.09, 1.66] 0.08 -0.16 [-0.76, 0.14] 0.39
@@
Table S15: Our curated taxonomy of manipulative strategies. We included an additional category, namely Others, to represent any responses that did not fit in the below categories, mainly in cases where manipulation was not required (e.g., general inquiries about logistics, affirmation, and reassurance). In addition, we excluded strategies that were irrelevant to the user-assistant setting (e.g., victimizing oneself, ridicule, and silent treatment) or suggested coercion (i.e., excessive use of force and authority).
@@
Others Responses to user inquiries where manipulation is not required "The price of this item is 60\$."
@@
Figure S9: The interaction page of our platform. On this page, participants were tasked to chat with the assigned AI agent for at least $10$ turns. After entering this page, the initial ratings were reset to the original value (=1=1), and none of the ratings were shared with the AI assistant, mimicking real-world human-AI interactions. Each option features a drop-down button (“See Details”) that displays its description. These descriptions were shown by default and could be hidden by pressing the arrow icon. Once the conversation surpassed the threshold of $10$ turns, the submit button became available, and participants could complete the task.